### v1

@@ 8. The local lifting argument
We used [formal-lifting-excellence-Gabber] to ensure the ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a $\mathbf{P}$-ring. A weaker result may be enough.
If lci implies $\mathbf{P}$, then every lci ring is a $\mathbf{P}$-ring (cf. [avramov-ci, (5.4)]). The ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a quotient of a regular ring as $R^{\wedge}$ is, so it is a $\mathbf{P}$-ring, avoiding [formal-lifting-excellence-Gabber]. This is the case in our applications.
We needed to do this because our $\mathfrak{c}(-)$ is only defined on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$. In [Nishimura-semilocal-lifting, BI-semilocal-lifting], $\mathfrak{c}(-)$ is defined on the whole of $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}-1}$ (Remark 8.4), so this is unnecessary.
@@ 9. Extending $\mathbf{P}$-assignments
Let $\mathbf{P},\mathbf{Q}$, $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ be as in Definition 8.3. For an integer $d\geq 0$ let ${}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be the full subcategory of $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ of rings $A\in\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ that are complete local of dimension $d$ whose $\mathbf{P}$-locus is the punctured spectrum. Let $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be the disjoint union of all ${}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. In other words, the objects are $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ are complete local rings in $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ whose $\mathbf{P}$-locus is the punctured spectrum, and the morphisms are local $\mathbf{P}$-maps whose closed fiber has dimension 00.
@@
Let $\mathbf{P},\mathbf{Q}$, $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}},\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be as in Definitions 8.3 and 9.1. Assume that $\mathbf{P}$ implies $(S_{1})$. Then every strictly functorial $\mathbf{P}$-assignment $\mathfrak{c}(-)$ on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ extends uniquely to a strictly functorial $\mathbf{P}$-assignment $\mathfrak{c}(-)$ on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ in a way that $\mathfrak{c}(A)$ has no embedded prime divisors for all $A\in\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
In the next two sections, we will find assignments on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ for $\mathbf{Q}$ trivial and $\mathbf{P}$=“$(S_{1})$,” $\mathbf{Q}$=“$(S_{1})$” and $\mathbf{P}$=“$(S_{2})$,” and $\mathbf{Q}$=“‘$(S_{2})$” and $\mathbf{P}$=“$(S_{k})$$(k\geq 3)$, “Gorenstein,” and “lci.”
@@ 10. $(S_{k})$-, Cohen–Macaulay-, and Gorenstein-assignments
Let $\mathbf{P}$=“$(S_{1})$.” Then $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0}_{\mathfrak{m}_{A}}(A))$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}_{\mathbf{P}}$.
@@
It is clear that $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary and $\mathfrak{c}(-)$ is strictly functorial, as the closed fiber of all maps in $\mathcal{A}_{\mathbf{P}}$ have dimension 00. As an Artinian ring is $(S_{1})$ we see ${}^{0}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}=\emptyset$, so $\dim A>0$ and any $\mathfrak{m}_{A}$-primary ideal is nonzero. ∎
Discussion 10.2.
Let $\mathbf{Q}$=“$(S_{1})$” and $\mathbf{P}$=“$(S_{2})$.” For $A\in\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$, let $\Sigma_{1}$ be the set of primary components $Q$ of 00 so that $\dim(A/Q)=1$, and let $\Sigma_{2}$ be the set of primary components $Q$ of 00 so that $\dim(A/Q)>1$. Note that $A$ is $(S_{1})$ and not $(S_{2})$, so 00 has no embedded primes and $\dim A>1$, therefore primary components of 00 are uniquely determined and $\Sigma_{2}\neq\emptyset$. Let $A_{1}=A/\bigcap_{Q\in\Sigma_{1}}Q,A_{2}=A/\bigcap_{Q\in\Sigma_{2}}Q$. Let $U_{2}$ be the punctured spectrum of $A_{2}$ and let $A_{2}^{\prime}=\mathcal{O}(U_{2})$. As $\dim(A/Q)>1$ for all $Q\in\Sigma_{2}$, similar to Lemma 4.9 (cf. [Macaulay-Cesnavi, Lemma 2.11]) we have $A_{2}^{\prime}$ is finite over $A_{2}$. This gives a finite birational ring map $A\to A_{1}\times A^{\prime}_{2}$. Let $\mathfrak{c}(A)$ be the conductor of this map.
Let $U$ (resp. $U_{1}$) be the punctured spectrum of $A$ (resp. $A_{1}$). Then we have $U=U_{1}\sqcup U_{2}$. Therefore $\mathfrak{c}(A)$ is either $\mathfrak{m}_{A}$-primary or $A$. As $U$ is $(S_{2})$ we have $U_{2}$ is $(S_{2})$, so $A_{2}^{\prime}$ is $(S_{2})$ by [EGA4_2, Théorème 5.10.5]. In particular $A\neq A_{1}\times A_{2}^{\prime}$ as $A$ is local and not $(S_{2})$, so $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary.
@@
Let $\mathbf{Q}$=“$(S_{1})$” and $\mathbf{P}$=“$(S_{2})$.” Then $\mathfrak{c}(-)$ as in Discussion 10.2 is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}_{\mathbf{P}}^{\mathbf{Q}}$.
@@
We know $\varphi$ has $(S_{1})$ fibers and its closed fiber has dimension 00. For $Q\in\Sigma_{1}$, $B/QB$ is therefore $1$-dimensional and $(S_{1})$, so all prime divisors $\mathfrak{P}$ of $QB$ are such that $\dim(B/\mathfrak{P})=1$. If we can show for all $Q\in\Sigma_{2}$ and all prime divisors $\mathfrak{P}$ of $QB$ (which are automatically minimal), we have $\dim(B/\mathfrak{P})>1$, then it will follow that $A_{1}\otimes_{A}B=B_{1}$ and $A_{2}\otimes_{A}B=B_{2}$, so $A_{2}^{\prime}\otimes_{A}B=B_{2}^{\prime}$ and $\mathfrak{c}(A)B=\mathfrak{c}(B)$.
@@
of rings. The ring $A_{2}^{\prime}\otimes_{A}B$ is $(S_{2})$ (as $A_{2}^{\prime}$ and the fibers of $\varphi$ are) and universally catenary, hence locally equidimensional [EGA4_2, Corollaire 5.10.9]. Let $\mathfrak{N}^{\prime}\in\operatorname{Max}(A_{2}^{\prime}\otimes_{A}B)$. Then $\mathfrak{N}^{\prime}\cap(A_{2}\otimes_{A}B)$ is the maximal ideal $\mathfrak{N}$ of the local ring $A_{2}\otimes_{A}B$, so $\mathfrak{N}^{\prime}\cap A_{2}$ is the maximal ideal of $A_{2}$, so $\mathfrak{N}^{\prime}\cap A_{2}^{\prime}\in\operatorname{Max}(A_{2}^{\prime})$. By flatness $\operatorname{ht}(\mathfrak{N}^{\prime})\geq\operatorname{ht}(\mathfrak{N}^{\prime}\cap A_{2}^{\prime})>1$. As $\mathfrak{N}^{\prime}$ was arbitrary, a similar discussion as the case of $A_{2}$ tells us for all $\mathfrak{Q}_{0}\in\operatorname{Min}(A_{2}\otimes_{A}B)$ we have $\dim((A_{2}\otimes_{A}B)/\mathfrak{Q}_{0})>1$, as desired. ∎
Discussion 10.4.
Let $A$ be an $(S_{2})$ Noetherian local ring that admits a normalized dualizing complex $\omega$. Then $A$ is catenary [stacks, Tag 0A80] and $(S_{2})$, so $A$ is equidimensional [EGA4_2, Corollaire 5.10.9]. We will use the standard facts [stacks, Tags 0A7U and 0A7V] of dualizing complexes without explicit reference.
@@
Discussion 10.5.
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“$(S_{k})$,” where $k\geq 3$. Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$, and let $p=0ptA$.
Let $\omega$ be a normalized dualizing complex of the complete local ring $A$. Let $\mathfrak{p}\in\operatorname{Spec}(A)$ be of height $d-1$. Then $\omega_{\mathfrak{p}}\in D^{\geq-d}(A_{\mathfrak{p}})$ and $H^{-d}(\omega_{\mathfrak{p}})\neq 0$. This tells us for all $b>-1-0ptA_{\mathfrak{p}}$, we have $H^{b}(\omega_{\mathfrak{p}})=0$. As $A_{\mathfrak{p}}$ is $(S_{k})$ we have $0ptA_{\mathfrak{p}}\geq\min\{d-1,k\}$, so for all $b>-1-\min\{d-1,k\}$, $H^{b}(\omega_{\mathfrak{p}})=0$. As $A$ is not $(S_{k})$, $p<\min\{d,k\}$, so $-p>-\min\{d,k\}\geq-1-\min\{d-1,k\}$. This tells us $H^{-p}(\omega_{\mathfrak{p}})=0$, in other words, the support of the nonzero module $H^{-p}(\omega)$ is $\{\mathfrak{m}_{A}\}$. By local and Matlis duality [stacks, Tags 0A84 and 08Z9] we see $H^{p}_{\mathfrak{m}_{A}}(A)$ is nonzero and of finite length.
@@
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“$(S_{k})$$(k\geq 3)$. Then $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0ptA}_{\mathfrak{m}_{A}}(A))$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
By Discussion 10.5 $H^{0ptA}_{\mathfrak{m}_{A}}(A)$ is of finite length, so $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary. Again, as an Artinian ring is $(S_{k})$ we see $\mathfrak{c}(A)\neq 0$. For all $\varphi:A\to B$ in $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$, the closed fiber of $\varphi$ has dimension 00, so we have $0ptA=0ptB$ [stacks, Tag 0337]. This shows $\mathfrak{c}(A)B=\mathfrak{c}(B)$. ∎
@@
It follows formally that for $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“Cohen–Macaulay” we have a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. Indeed, ${}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}=\emptyset$ for $d\leq 2$, and for $d>2$ and $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$ we let $\mathfrak{c}(A)$ be as in Lemma 10.6 for $k=d$. It also happens that for all $d$, the formula for $\mathfrak{c}(A)$ is the same, $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0ptA}_{\mathfrak{m}_{A}}(A))$.
Basics about Fitting ideals of a finite module can be found in [stacks, Tag 07Z6] and [Eisenbud-CA, Chapter 20]. The Fitting invariant of a finite module $M$ over a Noetherian ring $A$ is the first nonzero Fitting ideal of $M$. $M$ is projective of constant rank if and only if the Fitting invariant of $M$ is $A$, see [stacks, Tag 07ZD].
Discussion 10.8.
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“Gorenstein.” Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$.
Let $\omega$ be a normalized dualizing complex of the complete local ring $A$. When $d=0$, we let $\mathfrak{c}(A)$ be the Fitting invariant of the module $H^{0}(\omega)$, which is nonzero by definition, and is $\mathfrak{m}_{A}$-primary as $\dim A=0$ and as $\omega$ is not free. When $d>0$, let $\mathfrak{c}(A)=\operatorname{Fit}_{1}(H^{-d}(\omega))\cap\operatorname{Ann}_{A}(H^{1-d}(\omega))\cap\ldots\cap\operatorname{Ann}_{A}(H^{0}(\omega))$. As the punctured spectrum of $A$ is Gorenstein and as $\operatorname{Supp}(H^{-d}(\omega))=\operatorname{Spec}(A)$ (Discussion 10.4), we see $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary, and therefore nonzero as $\dim A>0$.
@@
Let $\mathbf{Q}$=“$(S_{2})$” and $\mathbf{P}$=“Gorenstein.” Then $\mathfrak{c}(-)$ as in Discussion 10.8 is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
We have seen $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary and nonzero. Strict functoriality follows immediately from the fact Fitting ideals commute with base change [stacks, Tag 07ZA] and that for $\varphi:A\to B$ in $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ and a normalized dualizing complex $\omega$ of $A$, $\omega\otimes^{L}_{A}B$ is a normalized dualizing complex of $B$. See for example [Lyu-dual-complex-lift, Lemma 7.1], note $\dim A=\dim B$. ∎

### v2

@@ 9. The local lifting argument
We used [KS21] to ensure the ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a $\mathbf{P}$-ring. A weaker result may be enough.
If lci implies $\mathbf{P}$, then every lci ring is a $\mathbf{P}$-ring (cf. [Avr99, (5.4)]). The ring $((R^{\wedge})_{\mathfrak{p}})^{*}$ is a quotient of a regular ring as $R^{\wedge}$ is, so it is a $\mathbf{P}$-ring, avoiding [KS21]. This is the case in our applications.
We needed to do this because our $\mathfrak{c}(-)$ is only defined on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$. In [Nis81, BI84], $\mathfrak{c}(-)$ is defined on the whole of $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}-1}$ (Remark 9.4), so this is unnecessary.
@@ 10. Extending $\mathbf{P}$-assignments
Let $\mathbf{P},\mathbf{Q}$, $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ be as in Definition 9.3. For an integer $d\geq 0$ let $^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be the full subcategory of $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ of rings $A\in\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ that are complete local of dimension $d$ whose $\mathbf{P}$-locus is the punctured spectrum. Let $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be the disjoint union of all $^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. In other words, the objects are $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ are complete local rings in $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ whose $\mathbf{P}$-locus is the punctured spectrum, and the morphisms are local $\mathbf{P}$-maps whose closed fiber has dimension 0.
@@
Let $\mathbf{P},\mathbf{Q}$, $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}},\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ be as in Definitions 9.3 and 10.1. Assume that $\mathbf{P}$ implies $(S_{1})$. Then every strictly functorial $\mathbf{P}$-assignment $\mathfrak{c}(-)$ on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ extends uniquely to a strictly functorial $\mathbf{P}$-assignment $\mathfrak{c}(-)$ on $\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$ in a way that $\mathfrak{c}(A)$ has no embedded prime divisors for all $A\in\mathcal{D}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
In the next two sections, we will find assignments on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ for $\mathbf{Q}$ trivial and $\mathbf{P}$$(S_{1})$,” $\mathbf{Q}$$(S_{1})$” and $\mathbf{P}$$(S_{2})$,” and $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$$(S_{k})$$(k\geq 3)$, “Gorenstein,” and “lci.”
@@ 11. $(S_{k})$-, Cohen–Macaulay-, and Gorenstein-assignments
Let $\mathbf{P}$$(S_{1})$.” Then $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0}_{\mathfrak{m}_{A}}(A))$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}_{\mathbf{P}}$.
@@
It is clear that $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary and $\mathfrak{c}(-)$ is strictly functorial, as the closed fiber of all maps in $\mathcal{A}_{\mathbf{P}}$ have dimension 0. As an Artinian ring is $(S_{1})$ we see $^{0}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}=\emptyset$, so $\dim A>0$ and any $\mathfrak{m}_{A}$-primary ideal is nonzero. ∎
Discussion 11.2.
Let $\mathbf{Q}$$(S_{1})$” and $\mathbf{P}$$(S_{2})$.” For $A\in\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$, let $\Sigma_{1}$ be the set of primary components $Q$ of 0 so that $\dim(A/Q)=1$, and let $\Sigma_{2}$ be the set of primary components $Q$ of 0 so that $\dim(A/Q)>1$. Note that $A$ is $(S_{1})$ and not $(S_{2})$, so 0 has no embedded primes and $\dim A>1$, therefore primary components of 0 are uniquely determined and $\Sigma_{2}\neq\emptyset$. Let $A_{1}=A/\bigcap_{Q\in\Sigma_{1}}Q,A_{2}=A/\bigcap_{Q\in\Sigma_{2}}Q$. Let $U_{2}$ be the punctured spectrum of $A_{2}$ and let $A_{2}^{\prime}=\mathcal{O}(U_{2})$. As $\dim(A/Q)>1$ for all $Q\in\Sigma_{2}$, similar to Lemma 4.9 (cf. [Čes21, Lemma 2.11]) we have $A_{2}^{\prime}$ is finite over $A_{2}$. This gives a finite birational ring map $A\to A_{1}\times A^{\prime}_{2}$. Let $\mathfrak{c}(A)$ be the conductor of this map.
Let $U$ (resp. $U_{1}$) be the punctured spectrum of $A$ (resp. $A_{1}$). Then we have $U=U_{1}\sqcup U_{2}$. Therefore $\mathfrak{c}(A)$ is either $\mathfrak{m}_{A}$-primary or $A$. As $U$ is $(S_{2})$ we have $U_{2}$ is $(S_{2})$, so $A_{2}^{\prime}$ is $(S_{2})$ by [EGA ̵IV2, Théorème 5.10.5]. In particular $A\neq A_{1}\times A_{2}^{\prime}$ as $A$ is local and not $(S_{2})$, so $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary.
@@
Let $\mathbf{Q}$$(S_{1})$” and $\mathbf{P}$$(S_{2})$.” Then $\mathfrak{c}(-)$ as in Discussion 11.2 is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}_{\mathbf{P}}^{\mathbf{Q}}$.
@@
We know $\varphi$ has $(S_{1})$ fibers and its closed fiber has dimension 0. For $Q\in\Sigma_{1}$, $B/QB$ is therefore $1$-dimensional and $(S_{1})$, so all prime divisors $\mathfrak{P}$ of $QB$ are such that $\dim(B/\mathfrak{P})=1$. If we can show for all $Q\in\Sigma_{2}$ and all prime divisors $\mathfrak{P}$ of $QB$ (which are automatically minimal), we have $\dim(B/\mathfrak{P})>1$, then it will follow that $A_{1}\otimes_{A}B=B_{1}$ and $A_{2}\otimes_{A}B=B_{2}$, so $A_{2}^{\prime}\otimes_{A}B=B_{2}^{\prime}$ and $\mathfrak{c}(A)B=\mathfrak{c}(B)$.
@@
of rings. The ring $A_{2}^{\prime}\otimes_{A}B$ is $(S_{2})$ (as $A_{2}^{\prime}$ and the fibers of $\varphi$ are) and universally catenary, hence locally equidimensional [EGA ̵IV2, Corollaire 5.10.9]. Let $\mathfrak{N}^{\prime}\in\operatorname{Max}(A_{2}^{\prime}\otimes_{A}B)$. Then $\mathfrak{N}^{\prime}\cap(A_{2}\otimes_{A}B)$ is the maximal ideal $\mathfrak{N}$ of the local ring $A_{2}\otimes_{A}B$, so $\mathfrak{N}^{\prime}\cap A_{2}$ is the maximal ideal of $A_{2}$, so $\mathfrak{N}^{\prime}\cap A_{2}^{\prime}\in\operatorname{Max}(A_{2}^{\prime})$. By flatness $\operatorname{ht}(\mathfrak{N}^{\prime})\geq\operatorname{ht}(\mathfrak{N}^{\prime}\cap A_{2}^{\prime})>1$. As $\mathfrak{N}^{\prime}$ was arbitrary, a similar discussion as the case of $A_{2}$ tells us for all $\mathfrak{Q}_{0}\in\operatorname{Min}(A_{2}\otimes_{A}B)$ we have $\dim((A_{2}\otimes_{A}B)/\mathfrak{Q}_{0})>1$, as desired. ∎
Discussion 11.4.
Let $A$ be an $(S_{2})$ Noetherian local ring that admits a normalized dualizing complex $\omega$. Then $A$ is catenary [Stacks, Tag 0A80] and $(S_{2})$, so $A$ is equidimensional [EGA ̵IV2, Corollaire 5.10.9]. We will use the standard facts [Stacks, Tags 0A7U and 0A7V] of dualizing complexes without explicit reference.
@@
Discussion 11.5.
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$$(S_{k})$,” where $k\geq 3$. Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$, and let $p=0ptA$.
Let $\omega$ be a normalized dualizing complex of the complete local ring $A$. Let $\mathfrak{p}\in\operatorname{Spec}(A)$ be of height d1d-1. Then $\omega_{\mathfrak{p}}\in D^{\geq-d}(A_{\mathfrak{p}})$ and $H^{-d}(\omega_{\mathfrak{p}})\neq 0$. This tells us for all $b>-1-0ptA_{\mathfrak{p}}$, we have $H^{b}(\omega_{\mathfrak{p}})=0$. As $A_{\mathfrak{p}}$ is $(S_{k})$ we have $0ptA_{\mathfrak{p}}\geq\min\{d-1,k\}$, so for all $b>-1-\min\{d-1,k\}$, $H^{b}(\omega_{\mathfrak{p}})=0$. As $A$ is not $(S_{k})$, $p<\min\{d,k\}$, so $p>-\min\{d,k\}\geq-1-\min\{d-1,k\}$. This tells us $H^{-p}(\omega_{\mathfrak{p}})=0$, in other words, the support of the nonzero module $H^{-p}(\omega)$ is $\{\mathfrak{m}_{A}\}$. By local and Matlis duality [Stacks, Tags 0A84 and 08Z9] we see $H^{p}_{\mathfrak{m}_{A}}(A)$ is nonzero and of finite length.
@@
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$$(S_{k})$$(k\geq 3)$. Then $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0ptA}_{\mathfrak{m}_{A}}(A))$ is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
By Discussion 11.5 $H^{0ptA}_{\mathfrak{m}_{A}}(A)$ is of finite length, so $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary. Again, as an Artinian ring is $(S_{k})$ we see $\mathfrak{c}(A)\neq 0$. For all $\varphi:A\to B$ in $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$, the closed fiber of $\varphi$ has dimension 0, so we have $0ptA=0ptB$ [Stacks, Tag 0337]. This shows $\mathfrak{c}(A)B=\mathfrak{c}(B)$. ∎
@@
It follows formally that for $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$=“Cohen–Macaulay” we have a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$. Indeed, $^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}=\emptyset$ for $d\leq 2$, and for $d>2$ and $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$ we let $\mathfrak{c}(A)$ be as in Lemma 11.6 for $k=d$. It also happens that for all $d$, the formula for $\mathfrak{c}(A)$ is the same, $\mathfrak{c}(A)=\operatorname{Ann}_{A}(H^{0ptA}_{\mathfrak{m}_{A}}(A))$.
Basics about Fitting ideals of a finite module can be found in [Stacks, Tag 07Z6] and [Eis95, Chapter 20]. The Fitting invariant of a finite module $M$ over a Noetherian ring $A$ is the first nonzero Fitting ideal of $M$. $M$ is projective of constant rank if and only if the Fitting invariant of $M$ is $A$, see [Stacks, Tag 07ZD].
Discussion 11.8.
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$=“Gorenstein.” Let $A\in{{}^{d}\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}}$.
Let $\omega$ be a normalized dualizing complex of the complete local ring $A$. When $d=0$, we let $\mathfrak{c}(A)$ be the Fitting invariant of the module $H^{0}(\omega)$, which is nonzero by definition, and is $\mathfrak{m}_{A}$-primary as $\dim A=0$ and as $\omega$ is not free. When $d>0$, let $\mathfrak{c}(A)=\operatorname{Fit}_{1}(H^{-d}(\omega))\cap\operatorname{Ann}_{A}(H^{1-d}(\omega))\cap\ldots\cap\operatorname{Ann}_{A}(H^{0}(\omega))$. As the punctured spectrum of $A$ is Gorenstein and as $\operatorname{Supp}(H^{-d}(\omega))=\operatorname{Spec}(A)$ (Discussion 11.4), we see $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary, and therefore nonzero as $\dim A>0$.
@@
Let $\mathbf{Q}$$(S_{2})$” and $\mathbf{P}$=“Gorenstein.” Then $\mathfrak{c}(-)$ as in Discussion 11.8 is a strictly functorial $\mathbf{P}$-assignment on $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$.
@@
We have seen $\mathfrak{c}(A)$ is $\mathfrak{m}_{A}$-primary and nonzero. Strict functoriality follows immediately from the fact Fitting ideals commute with base change [Stacks, Tag 07ZA] and that for $\varphi:A\to B$ in $\mathcal{A}^{\mathbf{Q}}_{\mathbf{P}}$ and a normalized dualizing complex $\omega$ of $A$, $\omega\otimes^{L}_{A}B$ is a normalized dualizing complex of $B$. See for example [Lyu, Lemma 7.1], note $\dim A=\dim B$. ∎